OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 3x + y = 1 6x + 2y = −9 0 2 There is no x value as there is no solution to this system x can be any value as there are infinitely many solutions to this system

OpenStudy (texaschic101):

do you prefer substitution or elimination ?

OpenStudy (anonymous):

Substitution.

OpenStudy (texaschic101):

3x + y = 1 --> y = -3x + 1 now sub -3x + 1 in for y in the 2nd equation Can you do that ?

OpenStudy (texaschic101):

Let me get you started... 6x + 2y = -9 6x + 2(-3x + 1) = -9 can you take it from here and solve for x ?

OpenStudy (anonymous):

What exactly do I do now?

OpenStudy (texaschic101):

you distribute the 2 through the parenthesis....can you do that ?

OpenStudy (anonymous):

I don't know what you mean.

OpenStudy (texaschic101):

let me show you... 6x + 2(-3x + 1) = -9 6x - 6x + 2 = -9 do you see what I did ?

OpenStudy (anonymous):

No.

OpenStudy (texaschic101):

ok.... 6x + 2(-3x + 1) = -9 we want to get rid of the parenthesis. To do this,we have to multiply the number right outside of the parenthesis, with everything in the parenthesis. 6x + 2 * -3x + 2*1 = -9 are you with me so far ?

OpenStudy (anonymous):

Yes.

OpenStudy (texaschic101):

then we get : 6x - 6x + 2 = -9 understanding ?

OpenStudy (anonymous):

Yeah.

OpenStudy (texaschic101):

we then want to move the 2 to the other side by subtracting 2 from both sides...like this 6x - 6x + 2 - 2 = -9 - 2 I subtracted 2 from both sides and now the equation reads : 6x - 6x = -11 do you understand ?

OpenStudy (anonymous):

Yeah.

OpenStudy (texaschic101):

6x - 6x = -11 0 = -11 when you have an equation where the end results do not equal, that means there is no solution. If the equation would have ended like this : 2 = 2 or 7 = 7, where it would equal, then there would be infinite solutions. Or if it would have ended like this: x = 4 or x = 6 then there would be one solution. So this system of equations has no solutions. Do you have any questions ?

OpenStudy (anonymous):

No questions. Thank you.

OpenStudy (texaschic101):

your welcome...glad I could help :)